### v1

@@ 3.2 Comparison to observations
These findings also agree well with our comparison of the observations to our hybrid fluid-particle jet model shown in Fig. 2, where we illustrate a ray-traced linearly polarised emission map of interpolated macro-particle attributes (a discussion of the used methodology is presented in the accompanying paper J. A. Kramer et al. in prep.). Specifically, we could accurately identify three of the four regions in our simulation: The 45∘ EVPA pattern on the south-western limb of our simulation is explained by the rotational viewing angle of about 35.5∘. If the jet would be viewed fully edge-on (see e.g. Kramer & MacDonald, 2021) they would be perpendicular to the jet direction. However, due to the rotation of the viewing angle, they appear as observed in region R of Fig. 1. The north-eastern limb of our observations (region R) is a feature that is also replicated with our simulations. In this case, it is predicted that the EVPAs should be perpendicular to the propagation direction, which is exactly what is observed with our VLBI data. On the spine (region R), the EVPAs are parallel to the jet propagation direction of the core region (Park et al., 2023), as predicted in a fully toroidal simulation (visible in Fig. 2 as well as in Kramer & MacDonald, 2021). The changing viewing angle might also explain the difference in total intensity observed between the two limbs (i.e. due to Doppler boosting). On the other hand, if this difference is an intrinsic feature of the 3C 84 jet, a highly twisted magnetic field (i.e. not purely toroidal) could still reconcile the observations with our simulations.
Region R is not present in our simulations, however, the polarisation enhancement there could be connected to the jet colliding with the interstellar medium and abruptly turning southwards scenario, proposed in Park et al. (2023). Through a kinematic and spectral study the authors showed that the jet could be changing its direction of propagation in that area. Specifically, they found that in that area the spectral index flattens out, hinting at the presence of a working surface of the radio jet. In that case, the magnetic field would be compressed there, resulting in an increase of its strength, manifested in a linear polarisation signal. Nevertheless, we note that our detection of the linearly polarised signal is marginal. This might indicate that, alternatively, the bend of the jet is intrinsic and the spectral inversion seen in Park et al. (2023) is due to opacity effects of the jet bend geometry seen at an oblique angle. Variability might also play a role, as newly ejected jet components might have altered the core region morphology. Our follow-up, high sensitivity EVN observations might be able to provide a more definitive conclusion.
@@ 3.3 Counter-jet depolarisation
Parallel to the signatures of polarised light detected in the core region of 3C 84, it is also of interest to note the absence of such signatures in the counter-jet region, as seen in Fig. 1. Here we assume the counter-jet to correspond to the diffuse emission seen North of the core region. Weaker polarisation signals in the receding jet side have already been reported (e.g. Laing, 1988) and Faraday depolarisation due to the presence of free-free absorbing circumnuclear gas (e.g. Walker et al., 2000; Fujita & Nagai, 2017; Wajima et al., 2020; Kino et al., 2018, 2021; Savolainen et al., 2023; Park et al., 2023, discuss the ambient medium around 3C 84) as one possible explanation. Using a noise estimated upper limit of for the linearly polarised flux of the counter-jet, we were able to compute the jet-to-counter-jet ratio of polarised light to be (under the assumption the approaching and receding jet sides have intrinsically the same brightness). This corresponds to an optically thick medium, characterised by an optical depth of (Rybicki & Lightman, 1979), which obscures the counter-jet. For comparison, we note that Wajima et al. (2020) measured higher values at 43 GHz and 86 GHz, using the Korean VLBI Network (KVN), but they were computing the total intensity jet-to-counter-jet ratio. Observations analysing the atomic and molecular gas content closing in on the central engine of 3C 84, similar to Nagai et al. (2019); Morganti et al. (2023); Oosterloo et al. (2024) will shed more light on the nature of the obscuring medium.
@@ 4 Conclusions
In this work we compared high-fidelity, high-resolution observations of the core region of 3C 84 in total intensity and linear polarisation to RMHD simulations. Our findings are summarised as follows:
@@
Linearly polarised flux is also detected in the region between these two limbs; the associated EVPAs exhibit a bimodal structure. They follow the bulk jet flow between the limbs but are perpendicularly oriented to it at the two limbs.
@@
Overall, our study presents an exploration of how state-of-the-art RMHD simulations can be utilised to gain new insights into jet launching in nearby AGN. In future studies, facilitated by the ever-increasing sensitivity of global VLBI arrays, such as the EVN array, additional radio sources will be scrutinised in more detail. In the same vein, our upcoming, multi-frequency follow-up observations of 3C 84 will shed more light on its core region, by utilising also spectral information.
@@ Acknowledgements.
- EHTC et al. (2022) EHTC, Akiyama, K., Alberdi, A., et al. 2022, ApJ, 930, L14
- EHTC et al. (2019) EHTC, Akiyama, K., Alberdi, A., et al. 2019, ApJ, 875, L4
@@ A.1 EVN 22 GHz data
3C 84 was observed with the EVN, augmented by additional antennas forming a global array, at 22 GHz in dual circular polarisation in November 2021, along with the calibrator sources NRAO 150 and J0102+58. The sources were observed by the following antennas: Effelsberg (EB), Jodrell Bank (JB), Onsala (ON), Metsähovi (MH), Medicina (MC), Yebes (YS), Brewster (BR), Fort Davis (FD), Kitt Peak (KP), Los Alamos (LA), Mauna Kea (MK), North Liberty (NL), Owens Valley (OV), Pie Town (PT), Hancock (HN), Saint Croix (SC), KVN Tamna (KT), KVN Ulsan (KU), KVN Yonsei (KY), Svetloe (SV), Urumqi (UR), and Zelenchukskaya (ZC) in full track mode for a total of 26 h. The data were recorded at an aggregate bitrate of 2 Gbps with two-bit digitisation. The resulting -coverage for 3C 84 is shown in Fig. 3. The data were then channelised in eight intermediate frequencies (IFs) per polarisation, with an IF bandwidth of 32 MHz; the data correlation was performed in zoom band mode resulting in a total recording bandwidth of MHz. Finally, the data were correlated using the SFXC correlator (Keimpema et al. 2015) at Joint Institute for Very Long Baseline Interferometry European Research Infrastructure Consortium (JIVE).
@@ A.2 Calibration and imaging
The next step in the data evaluation involved the post-correlation and calibration steps. We used the rPICARD pipeline (Janssen et al. 2019) to perform the standard VLBI fringe-fitting and calibration procedures. In short, we calibrated the phases by removing the delay and phase offsets between the IFs and then applied a global fringe fitting algorithm over the full bandwidth. The visibility amplitudes were calibrated using the measured system noise temperature; the gain curve was also applied for each station. For the stations that do not take atmospheric opacity effects into account, we applied the necessary corrections with the inferred receiver temperature and opacity along the line of sight. Since we were interested in the polarised emission, we also removed the cross-hand phase and delay offsets. More details about the VLBI calibration can be found, for example, in Martí-Vidal et al. (2012); Hada et al. (2016); Kim et al. (2019).
@@
The data were exported and then frequency-averaged and time-averaged (coherently, in thirty second bins) to be imaged using the difmap software (Shepherd et al. 1994). Outliers in the visibility domain were flagged. A uniform weighting scheme was used in the images presented here, with the goal of sufficiently resolving the fine, linearly polarised structures in the core. Naturally weighted images were also produced to search for more extended linearly polarised signal in the outer regions; however, their S/N is not sufficient for a robust detection and we therefore refrain from presenting them here. We then imaged the sources in total intensity, using the standard, iterative procedure of CLEANing and self-calibration in phase and amplitude. Finally, the total flux was scaled up to of the total single dish flux measurements, taken with EF before, during, and after the observations. The compactness factor was estimated based on a comparison between the 15 GHz variability light curve (Paraschos et al. 2023) and the corresponding VLBI flux of 3C 84 (Paraschos et al. 2022, compare also with the MOJAVE333https://www.cv.nrao.edu/MOJAVE/ database Lister et al. 2018). The uncertainty in the averaged single dish measurement is of the order of 0.5%.
@@
In Fig. 4 we show a zoomed-in version of the limb brightening of 3C 84 in Stokes . In order to better examine this high region, we reconstructed the image with the software package ehtim (Chael et al. 2016, 2018). Among the forward modelling approaches, the regularised maximum likelihood (RML) method is an image synthesis family that require maximising the likelihood of a set of visibility data, given a set of predicted model visibility values and regularisers. RML imaging differs from traditional methods like CLEAN by offering a non-parametric approach to reconstructing images directly from the data, achieving better angular resolution and image fidelity (e.g. EHTC et al. 2019, 2022; Fuentes et al. 2023).
@@
For the first imaging iteration, we used a prior image of a Gaussian model with a size of 250 as FWHM, across a 6.5 mas field of view, and mapped on a pixel grid. We incorporated closure phases and logarithmic closure amplitudes into the imaging process, accounting for non-closing systematic errors of 1%. Having achieved convergence, we proceeded with self-calibration. In the second imaging run, we included visibility phases and amplitudes in the process. The third and last imaging round we included both closure quantities and complex visibilities. To ensure the correct choice of relative weights for the regularisation terms, we conducted a parameter survey, testing different values for the relative entropy parameter (MEM), total variation (tv), total squared variation (tv2), and the 1 norm (see also Savolainen et al. 2023). The final combination of values was chosen based on the criterion that provided the best fit to the data, producing the highest fidelity image are relative MEM: 0.1, tv: 0.3, tv2: 0, and 1: 0.1.
@@ A.3 Polarisation
Determining the absolute positions of the EVPAs requires prior knowledge of quasi-simultaneous single dish, polarised flux measurements, or an estimate of how the D-terms change with time. At 22 GHz no such measurements were available to us. We therefore assumed that the EVPAs in the downstream jet area of the calibrator NRAO 150 are intrinsically aligned to the bulk jet flow, as traced by the ridge line444See Paraschos et al. (2022) for details about the ridge line determination. (see, e.g. Pushkarev et al. 2005; Hada et al. 2016). Figure 5 displays NRAO 150 in Stokes and polarised flux density. We note that, as shown in Molina et al. (2014) for past observations at 22 GHz, some EVPA orientation variability with time is present in the core region. In the downstream, optically thin region, however, the detections are marginal and only in some epochs, so they do not allow for an estimate of their temporal variability. The calculated correction is . We also note with interest the filamentary structure in linear polarisation present in this source (see Fig. 5), which we will investigate in an upcoming work (Debbrecht et al. in prep.).
Since inaccurately determined D-terms could affect the revealed spine-sheath polarisation pattern, we sought to determine them in an alternative way. We used a pipeline based on the DoG-HiT software (Müller & Lobanov 2022); this approach has the added benefit of also quantifying the uncertainty in the absolute position of the EVPAs. The result is displayed in Fig. 6. The EVPA patterns are in agreement with the fiducial image reconstruction shown in Fig. 1. Region R of the jet bend is recovered, as well as the jet sheath (regions R and R). Region R, which in our interpretation corresponds to the spine, is blended together with R here. A comparison between the EVPAs in the two implementations yields a total uncertainty of , while the average difference in corrections from the D-terms is of the order of . Our approach discussed above is based on a number of assumptions, yielding uncertainties in the presented absolute EVPA orientation. We note, however, that the EVPAs in the sheath and spine are at right angles, regardless of the correction factor applied to determine their absolute orientation.
@@ B.1 Setup
In the PLUTO Code (Mignone et al. 2007), the framework includes the equations of special RMHD, where a set of equations represents the vector of conservative variables and their associated fluxes, that is,
@@
These simulations operate with dimensionless grid units, requiring the conversion of the RMHD jet flow’s thermal properties into physical units during post-processing. Using dimensionless quantities helps avoid extremely small or large numbers in the simulation. Nonetheless, to achieve an accurate representation, physical scaling to include specific scales of length, time, and energy is essential (see, e.g. Mignone et al. 2007; Kramer & MacDonald 2021). RMHD simulations struggle to capture the kinetic-scale physics necessary for producing the non-thermal electron distributions behind observed synchrotron emissions. However, our recent jet simulations (see Fig. 2 and J. A. Kramer et al. in prep. for details), suitable for parsec-scale jets, incorporate microphysics to automatically create power-law electron distributions. Vaidya et al. (2018) developed a novel hybrid framework for particle transport within the PLUTO code555For details refer to http://plutocode.ph.unito.it. (Mignone et al. 2007) that enhances the understanding of high-energy non-thermal emissions using Lagrangian particles from 3D RMHD simulations. This approach leverages the simulation’s magnetic field to assess radiative losses, including synchrotron cooling, and integrates micro-physical aspects of spectral evolution based on fluid dynamics. The inclusion of Lagrangian particles in these simulations has proven instrumental in revealing detailed fluid characteristics such as velocity, density, and magnetic field strength, enabling an in-depth analysis of plasma transport, magnetic field impacts, and astrophysical phenomena like particle acceleration and shock dynamics. This methodology provides a more nuanced view of complex processes in astrophysical jets, aligning closely with observational data (Kramer 2023).
@@
RADMC-3D, a robust ray-tracing code developed by Dullemond et al. (2012) for astrophysical radiative transfer issues, calculates the Stokes parameters , , , for total, linear, and circular polarisation. It handles synchrotron absorption, emissivity, Faraday conversion, and rotation (MacDonald & Marscher 2018). The output, influenced by optical depth, normalisation constant, power-law distribution’s low-energy cutoff, magnetic field strength, and orientation, offers a detailed view of polarised synchrotron radiation. Unlike Kramer & MacDonald (2021), we use MacDonald & Nishikawa (2021) enhanced Stokes module in RADMC-3D, introducing a full rotational perspective to our 3D solution and adapting it to account for variable spectral indices. Finally, we performed a similar procedure for scaling the simulated flux densities to the observed ones, as discussed in Appendix A.2. We note that the simulated fractional linear polarisation is higher than that of the observations because in our simulations depolarising agents, such as foreground Faraday screens and free-free absorbing media, were not taken into consideration.

### v2

@@ 3.2 Comparison to observations
These findings also agree well with our comparison of the observations to our hybrid fluid-particle jet model shown in Fig. 2, where we illustrate a ray-traced linearly polarised emission map of interpolated macro-particle attributes (a discussion of the used methodology is presented in the accompanying paper J. A. Kramer et al. in prep.). Specifically, we could accurately identify three of the four regions in our simulation. The 45∘ EVPA pattern on the south-western limb of our simulation is explained by the rotational viewing angle of about 35.5∘. If the jet were viewed fully edge-on (see e.g. Kramer & MacDonald, 2021) the EVPA pattern would be perpendicular to the jet direction. However, due to the rotation of the viewing angle, they appear as observed in region R of Fig. 1. The north-eastern limb of our observations (region R) is a feature that is also replicated with our simulations. In this case, it is predicted that the EVPAs should be perpendicular to the propagation direction, which is exactly what is observed with our VLBI data. On the spine (region R), the EVPAs are parallel to the jet propagation direction of the core region (Park et al., 2023), as predicted in a fully toroidal simulation (visible in Fig. 2 and in Kramer & MacDonald, 2021). The changing viewing angle might also explain the difference in total intensity observed between the two limbs (i.e. due to Doppler boosting). On the other hand, if this difference is an intrinsic feature of the 3C 84 jet, a highly twisted magnetic field (i.e. not purely toroidal) could still reconcile the observations with our simulations.
Region R is not present in our simulations; however, the polarisation enhancement there could be connected to the jet colliding with the interstellar medium and abruptly turning southwards scenario that was proposed in Park et al. (2023). Through a kinematic and spectral study the authors showed that the jet could be changing its direction of propagation in that area. Specifically, they found that in that area the spectral index flattens out, hinting at the presence of a working surface of the radio jet. In that case, the magnetic field would be compressed there, resulting in an increase in its strength, manifested in a linear polarisation signal. Nevertheless, we note that our detection of the linearly polarised signal is marginal. This might indicate that, alternatively, the bend of the jet is intrinsic and the spectral inversion seen in Park et al. (2023) is due to opacity effects of the jet bend geometry seen at an oblique angle. Variability might also play a role, as newly ejected jet components might have altered the core region morphology. Our follow-up high-sensitivity EVN observations might be able to provide a more definitive conclusion.
@@ 3.3 Counter-jet depolarisation
Parallel to the signatures of polarised light detected in the core region of 3C 84, it is also of interest to note the absence of such signatures in the counter-jet region, as seen in Fig. 1. Here we assume the counter-jet to correspond to the diffuse emission seen north of the core region. Weaker polarisation signals in the receding jet side have already been reported (e.g. Laing, 1988) and Faraday depolarisation due to the presence of free-free absorbing circumnuclear gas (see discussion of the ambient medium around 3C 84 in e.g. Walker et al., 2000; Fujita & Nagai, 2017; Wajima et al., 2020; Kino et al., 2018, 2021; Savolainen et al., 2023; Park et al., 2023) as one possible explanation. Using a noise estimated upper limit of for the linearly polarised flux of the counter-jet, we were able to compute the jet-to-counter-jet ratio of polarised light to be (under the assumption that the approaching and receding jet sides have intrinsically the same brightness). This corresponds to an optically thick medium, characterised by an optical depth of (Rybicki & Lightman, 1979), which obscures the counter-jet. For comparison, we note that Wajima et al. (2020) measured higher values at 43 GHz and 86 GHz, using the Korean VLBI Network (KVN), but they were computing the total intensity jet-to-counter-jet ratio. Observations analysing the atomic and molecular gas content closing in on the central engine of 3C 84, similar to those performed by Nagai et al. (2019), Morganti et al. (2023), and Oosterloo et al. (2024) will shed more light on the nature of the obscuring medium.
@@ 4 Conclusions
In this work we compared high-fidelity high-resolution observations of the core region of 3C 84 in total intensity and linear polarisation to RMHD simulations. Our findings are summarised as follows:
@@
Linearly polarised flux is also detected in the region between these two limbs; the associated EVPAs exhibit a bimodal structure. They follow the bulk jet flow between the limbs, but are perpendicularly oriented to it at the two limbs.
@@
Overall, our study presents an exploration of how state-of-the-art RMHD simulations can be utilised to gain new insights into jet launching in nearby AGN. In future studies, facilitated by the ever-increasing sensitivity of globe-spanning VLBI arrays, such as the EVN array, additional radio sources will be scrutinised in more detail. In the same vein, our upcoming, multi-frequency follow-up observations of 3C 84 will shed more light on its core region by also utilising spectral information.
@@ Acknowledgements.
- EHTC et al. (2024) EHTC, Akiyama, K., Alberdi, A., et al. 2024, A&A, 681, A79
- EHTC et al. (2022a) EHTC, Akiyama, K., Alberdi, A., et al. 2022a, ApJ, 930, L12
- EHTC et al. (2022b) EHTC, Akiyama, K., Alberdi, A., et al. 2022b, ApJ, 930, L14
- EHTC et al. (2019a) EHTC, Akiyama, K., Alberdi, A., et al. 2019a, ApJ, 875, L1
- EHTC et al. (2019b) EHTC, Akiyama, K., Alberdi, A., et al. 2019b, ApJ, 875, L4
@@ A.1 EVN 22 GHz data
The radio source 3C 84, along with the calibrator sources NRAO 150 and J0102+58, was observed with the EVN, augmented by additional antennas forming a globe-spanning array, at 22 GHz in dual circular polarisation in November 2021. The sources were observed by the following antennas: Effelsberg (EB), Jodrell Bank (JB), Onsala (ON), Metsähovi (MH), Medicina (MC), Yebes (YS), Brewster (BR), Fort Davis (FD), Kitt Peak (KP), Los Alamos (LA), Mauna Kea (MK), North Liberty (NL), Owens Valley (OV), Pie Town (PT), Hancock (HN), Saint Croix (SC), KVN Tamna (KT), KVN Ulsan (KU), KVN Yonsei (KY), Svetloe (SV), Urumqi (UR), and Zelenchukskaya (ZC) in full track mode for a total of 26 h. The data were recorded at an aggregate bitrate of 2 Gbps with two-bit digitisation. The resulting -coverage for 3C 84 is shown in Fig. 3. The data were then channelised in eight intermediate frequencies (IFs) per polarisation, with an IF bandwidth of 32 MHz; the data correlation was performed in zoom band mode resulting in a total recording bandwidth of MHz. Finally, the data were correlated using the SFXC correlator (Keimpema et al. 2015) at Joint Institute for Very Long Baseline Interferometry European Research Infrastructure Consortium (JIVE).
@@ A.2 Calibration and imaging
The next step in the data evaluation involved the post-correlation and calibration steps. We used the rPICARD pipeline (Janssen et al. 2019) to perform the standard VLBI fringe-fitting and calibration procedures. In short, we calibrated the phases by removing the delay and phase offsets between the IFs and then applied a global fringe fitting algorithm over the full bandwidth. The visibility amplitudes were calibrated using the measured system noise temperature; the gain curve was also applied for each station. For the stations that do not take atmospheric opacity effects into account, we applied the necessary corrections with the inferred receiver temperature and opacity along the line of sight. Since we were interested in the polarised emission, we also removed the cross-hand phase and delay offsets. More details about the VLBI calibration can be found, for example, in Martí-Vidal et al. (2012), Hada et al. (2016), and Kim et al. (2019).
@@
The data were exported and then frequency-averaged and time-averaged (coherently, in 30-second bins) to be imaged using the difmap software (Shepherd et al. 1994). Outliers in the visibility domain were flagged. A uniform weighting scheme was used in the images presented here, with the goal of sufficiently resolving the fine linearly polarised structures in the core. Naturally weighted images were also produced to search for more extended linearly polarised signal in the outer regions; however, their S/N was not sufficient for a robust detection and we therefore refrain from presenting them here. We then imaged the sources in total intensity, using the standard, iterative procedure of CLEANing and self-calibration in phase and amplitude. Finally, the total flux was scaled up to of the total single dish flux measurements, taken with EF before, during, and after the observations. The compactness factor was estimated based on a comparison between the 15 GHz variability light curve (Paraschos et al. 2023) and the corresponding VLBI flux of 3C 84 (Paraschos et al. 2022, compare also with the MOJAVE333https://www.cv.nrao.edu/MOJAVE/ database Lister et al. 2018). The uncertainty in the averaged single-dish measurement is of the order of 0.5%.
@@
In Fig. 4 we show a zoomed-in version of the limb brightening of 3C 84 in Stokes . In order to better examine this high region, we reconstructed the image with the software package ehtim (Chael et al. 2016, 2018). Among the forward modelling approaches, the regularised maximum likelihood (RML) method is an image synthesis family that require maximising the likelihood of a set of visibility data, given a set of predicted model visibility values and regularisers. RML imaging differs from traditional methods like CLEAN by offering a non-parametric approach to reconstructing images directly from the data, achieving better angular resolution and image fidelity (e.g. EHTC et al. 2019b, 2022b; Fuentes et al. 2023).
@@
For the first imaging iteration, we used a prior image of a Gaussian model with a size of 250 as of full width at half maximum, across a 6.5 mas field of view, and mapped on a pixel grid. We incorporated closure phases and logarithmic closure amplitudes into the imaging process, accounting for non-closing systematic errors of 1%. Having achieved convergence, we proceeded with self-calibration. In the second imaging run, we included visibility phases and amplitudes in the process. The third and last imaging round we included both closure quantities and complex visibilities. To ensure the correct choice of relative weights for the regularisation terms, we conducted a parameter survey, testing different values for the relative entropy parameter (MEM), total variation (tv), total squared variation (tv2), and the 1 norm (see also Savolainen et al. 2023). The final combination of values was chosen based on the criterion that provided the best fit to the data, producing the highest fidelity image at relative MEM: 0.1, tv: 0.3, tv2: 0, and 1: 0.1.
@@ A.3 Polarisation
Determining the absolute positions of the EVPAs requires prior knowledge of quasi-simultaneous single-dish polarised flux measurements or an estimate of how the D-terms change with time. At 22 GHz no such measurements were available to us. We therefore assumed that the EVPAs in the downstream jet area of the calibrator NRAO 150 are intrinsically aligned to the bulk jet flow, as traced by the ridge line444See Paraschos et al. (2022) for details about the ridge line determination. (see e.g. Pushkarev et al. 2005; Hada et al. 2016). Figure 5 displays NRAO 150 in Stokes and polarised flux density. As shown in Molina et al. (2014) for past observations at 22 GHz, we note that some EVPA orientation variability with time is present in the core region. In the downstream optically thin region, however, the detections are marginal and only in some epochs, so they do not allow an estimate of their temporal variability. The calculated correction is . We also note with interest the filamentary structure in linear polarisation present in this source (see Fig. 5), which we will investigate in an upcoming work (Debbrecht et al. in prep.).
Since inaccurately determined D-terms could affect the revealed spine-sheath polarisation pattern, we sought to determine them in an alternative way. We used a pipeline based on the DoG-HiT software (Müller & Lobanov 2022); this approach has the added benefit of also quantifying the uncertainty in the absolute position of the EVPAs. The result is displayed in Fig. 6. The EVPA patterns are in agreement with the fiducial image reconstruction shown in Fig. 1. Region R of the jet bend is recovered, as is the jet sheath (regions R and R). Region R, which in our interpretation corresponds to the spine, is blended together with R here. A comparison between the EVPAs in the two implementations yields a total uncertainty of , while the average difference in corrections from the D-terms is of the order of . Our approach discussed above is based on a number of assumptions, yielding uncertainties in the presented absolute EVPA orientation. We note, however, that the EVPAs in the sheath and spine are at right angles, regardless of the correction factor applied to determine their absolute orientation.
@@ B.1 Set-up
In the PLUTO Code (Mignone et al. 2007) the framework includes the equations of special RMHD, where a set of equations represents the vector of conservative variables and their associated fluxes:
@@
These simulations operate with dimensionless grid units, requiring the conversion of the RMHD jet flow’s thermal properties into physical units during post-processing. Using dimensionless quantities helps avoid extremely small or large numbers in the simulation. Nonetheless, to achieve an accurate representation, physical scaling to include specific scales of length, time, and energy is essential (see e.g. Mignone et al. 2007; Kramer & MacDonald 2021). RMHD simulations struggle to capture the kinetic-scale physics necessary for producing the non-thermal electron distributions behind observed synchrotron emissions. However, our recent jet simulations (see Fig. 2 and J. A. Kramer et al. in prep. for details), suitable for parsec-scale jets, incorporate microphysics to automatically create power-law electron distributions. Vaidya et al. (2018) developed a novel hybrid framework for particle transport within the PLUTO code555For details refer to http://plutocode.ph.unito.it. (Mignone et al. 2007) that enhances the understanding of high-energy non-thermal emissions using Lagrangian particles from 3D RMHD simulations. This approach leverages the simulation’s magnetic field to assess radiative losses, including synchrotron cooling, and integrates micro-physical aspects of spectral evolution based on fluid dynamics. The inclusion of Lagrangian particles in these simulations has proven instrumental in revealing detailed fluid characteristics such as velocity, density, and magnetic field strength, enabling an in-depth analysis of plasma transport, magnetic field impacts, and astrophysical phenomena like particle acceleration and shock dynamics. This methodology provides a more nuanced view of complex processes in astrophysical jets, aligning closely with observational data (Kramer 2023).
@@
The robust ray-tracing code RADMC-3D, developed by Dullemond et al. (2012) for astrophysical radiative transfer issues, calculates the Stokes parameters , , , for total, linear, and circular polarisation. It handles synchrotron absorption, emissivity, Faraday conversion, and rotation (MacDonald & Marscher 2018). The output, influenced by optical depth, normalisation constant, power-law distribution’s low-energy cutoff, magnetic field strength, and orientation, offers a detailed view of polarised synchrotron radiation. Unlike Kramer & MacDonald (2021), we used the enhanced Stokes module in RADMC-3D (MacDonald & Nishikawa 2021), introducing a full rotational perspective to our 3D solution and adapting it to account for variable spectral indices. Finally, we performed a similar procedure for scaling the simulated flux densities to the observed ones, as discussed in Appendix A.2. We note that the simulated fractional linear polarisation is higher than that of the observations because in our simulations depolarising agents, such as foreground Faraday screens and free-free absorbing media, were not taken into consideration.